OpenStudy (anonymous):

Simplify the given equation. 5x + 2(x - 3) = -2(x - 1) 7x - 6 = -2x - 2 7x - 6 = -2x + 2 7x - 3 = -2x - 1

OpenStudy (anonymous):

its easyy

OpenStudy (anonymous):

Use distributive... answer is 7x-6=-2x+2

Parth (parthkohli):

=> 5x + 2x - 6 = -2x + 2 = 7x - 6 = -2x + 2 = 9x = 8 = x = 8/9 Solve all the ones left by transposing. Please try some on your own as you have to learn how to do it.

OpenStudy (anonymous):

need work??

OpenStudy (anonymous):

but if you simplify by adding the like terms together,you can can easily find out the answer...

OpenStudy (anonymous):

it would be better if you post the way you how you worked it out.... than getting the answer as ultimately you and you alone have to solve this in your exam...then no none is gonna help you...hence try yourself first....

OpenStudy (anonymous):

5x+2(x-3)=-2(x-1) 5x+2x-6=-2x+2 7x-6=-2x+2 7x+2x=2+6 9x=8 x=8/9